# Dynamics of Cosmic Superstrings and the Overshoot Problem

### v1

@@ IV Oscillating tension
$T_{\phi}\gg T_{R}\quad\Leftrightarrow\quad\epsilon m\ll 1\,$, (51)
one can approximate $a^{2}\dot{R}^{2}$ by its average $\langle a^{2}\dot{R}^{2}\rangle=1/2$, integrate (5) and reduce the dynamics to:
$\ddot{R}+\left(2H+\frac{\dot{\mu}}{2\mu}\right)\dot{R}\,+\frac{R}{\epsilon_{0}^{2}}\frac{\mu}{\mu_{0}}=0\,.$ (52)
@@
$\ddot{\sigma}+\left(\frac{\mu}{\mu_{0}\epsilon_{0}^{2}}+\frac{3}{16}\frac{\dot{\mu}^{2}}{\mu^{2}}-\frac{1}{4}\frac{\ddot{\mu}}{\mu}-\frac{\ddot{a}}{a}-\frac{1}{2}\frac{\dot{a}\dot{\mu}}{a\mu}\right)\sigma=0\,.$ (54)
Noting that when the modulus oscillates around its minimum, so does the string tension, (54) is a Hill equation. The oscillating tension may open up a rich dynamics for the circular strings, in particular the possibility of resonances.
As is often the analysis of Mathieu/Hill’s equations, it is convenient to rescale the time variable such that the period of the field oscillation is exactly $\pi$. We therefore define $\tau\equiv\frac{m}{2}t$ such that (54) becomes:
$\sigma^{\prime\prime}+\left(4\frac{\mu/\mu_{0}}{m^{2}\epsilon_{0}^{2}}+\frac{3}{16}\frac{\mu^{\prime 2}}{\mu^{2}}-\frac{1}{4}\frac{\mu^{\prime\prime}}{\mu}-\frac{a^{\prime\prime}}{a}-\frac{1}{2}\frac{a^{\prime}\mu^{\prime}}{a\mu}\right)\sigma=0\,.$ (55)
Equation (55) is an accurate description of the circular string dynamics, governed by (5) and (6), in the regime $m\epsilon_{0}\ll 1$. In this limit one may approximate (55) by keeping only the first term inside the brackets, allowing for an approximate WKB solution of the form:
$\sigma(\tau)=\sigma_{0}\left(\frac{\mu_{0}}{\mu(\tau)}\right)^{1/4}\cos\left[\frac{2}{m\epsilon_{0}}\int_{0}^{\tau}d\tilde{\tau}\left(\frac{\mu_{0}}{\mu(\tilde{\tau})}\right)^{1/2}+\delta\right],$ (56)
where $\delta$ is a phase fixed by the initial conditions. We see that the amplitude of the oscillations is modulated by the time variation of the tension. We stress that (56) holds both in flat space and in the presence of gravity. Note also that this result is independent of whether the variation of the string tension is monotonic (as in Conlon et al. (2024)) or oscillating, provided it varies sufficiently slowly, cf. (51). In the WKB regime the physical radius of the string scales as $R_{\rm max}\propto a^{-1}\mu^{-1/2}$, and so it grows if:
$\frac{d}{dt}(a\mu^{1/2})<0\Leftrightarrow\frac{\dot{\mu}}{\mu}<-2H\,$, (57)
that is, if the growth condition of Conlon et al. (2024) is respected.
IV.1 Flat spacetime
In order to build some physical intuition for the behaviour of the strings with an oscillating tension, let us first analyse the flat space solutions of (5) and (6). In flat space $a={\rm const}$, $H=0$ and the scalar motion is harmonic as there is no friction:
$\phi=\phi_{0}\cos(mt)=\phi_{0}\cos(2\tau)\,.$ (58)
@@ IV.1 Flat spacetime
$\mu=\mu_{0}\ e^{-\sqrt{6}\beta\,\phi_{0}\cos(2\tau)}\,.$ (59)
From (53) and (56) one then sees that the radius of the string follows:
$R(\tau)\propto\mu(\tau)^{-1/2}\,$, (60)
@@
Having seen that in the WKB regime the system presents no instability, we adapt the standard methods of Floquet’s analysis (see e.g. Amin et al. (2014)) to the current problem and look for instabilities when (51) is not satisfied. Defining the Floquet exponent $\mu_{F}\in\mathbb{R}$ via:
$R(\tau)=e^{\mu_{F}\,\tau}f(\tau)\,$, (61)
@@
Figure 12: No instability in flat space: growth exponent as a function of the timescale of the problem $\epsilon_{0}^{2}m^{2}$ (obtained from a numerical analysis of (5)-(6) with $\beta\phi_{0}=\sqrt{2/3}$.)
@@ IV.2 Curved spacetime
$H^{2}=\frac{1}{3M_{p}^{2}}\left(\frac{1}{2}\dot{\phi}^{2}+V(\phi)\right).$ (62)
As above, we are interested in a regime where the scalar is oscillating around the minimum of its potential. If the potential is well approximated by a quadratic term, the solutions to (50) and (62) are of the form:
@@
$\mu=\mu_{0}\ e^{-\sqrt{6}\beta\,\frac{\phi_{0}}{\tau}\cos(\tau)}\,.$ (64)
We see that in the late time limit, the variation of the tension becomes less significant. Taking into account this time dependence, one can rewrite the solution (56) in terms of the string radius as:
$R(\tau)\propto\frac{\mu(\tau)^{-1/2}}{a(\tau)}\,.$ (65)
@@ V Discussion and conclusions
The first part of this work has been devoted to studying whether the rolling volume mode actually settles in its minimum or overshoots it causing a decompactification limit. In fact, the typical LVS scalar potential for the canonically normalised volume mode $\Phi$ (9) features a minimum and a maximum. If $\Phi$ reaches the minimum with too much kinetic energy, the field will overshoot the maximum and run away to infinity. In the absence of any background fluid, this overshoot problem is present for generic initial conditions. The canonical way to overcome this problem is to introduce a fraction of radiation in the system. Depending on the specific embedding of the model, this solution may or may not be well justified. Therefore, following the work of Sánchez González et al. (2025), we studied how the inclusion of a string loop fluid in the dynamical system affects the overshooting problem.
@@
We first analysed the behaviour of the dynamical system including cosmic string loops in the absence of a radiation background. We performed a stability analysis of the fixed points, summarised in the stability map in Fig. 1. For an exponential potential with $\lambda=3$, as for the approximated LVS potential (26), we found that the final attractor of a system with only F-strings ($\beta=1/2$) is the loop tracker $\mathcal{L}$, confirming the result of Sánchez González et al. (2025). For $\beta=1/2$, the kinetic energy of the field is too high to stop at the minimum, and the system overshoots. On the other hand, for effective strings coming from D3-branes wrapping an internal 2-cycle ($\beta=1/3$) or NS5-branes wrapping a 4-cycle ($\beta=1/6$), the final attractor of the system would be the mixed tracker $\mathcal{T}_{1}$. As the convergence towards this fixed point is slow, the system still spends a long time hovering around $\mathcal{L}$, even if this critical point is unstable. Hence, whether the field overshoots or converges to the minimum is still dependent on the fraction of kinetic energy at $\mathcal{L}$. For $\beta=1/3$, the situation is qualitatively similar to the F-string case, with a kinetic energy that is too high at the minimum to avoid overshooting. For $\beta=1/6$, the NS5-string loops are so effective at absorbing energy from the field that they can indeed prevent it from overshooting. In fact, for NS5-strings, at $\mathcal{L}$ the kinetic energy of the field constitutes only a small fraction of the total energy density, $\Omega_{\rm k}=\beta^{2}\simeq 0.03$, while the fraction of energy density in NS5-string loops gets over $97\%$, as shown in Fig. 4. Moreover, when the field starts oscillating around the minimum, the tension of the strings stops decreasing. Therefore, the loops start redshifting as matter, just like the field. This means that the ratio of their energy densities stays constant throughout the period of early matter domination, provided that the strings do not lose energy through GW or particle emission. The most interesting aspect is that this fraction is always rather high, ranging from $25\%$ to $60\%$. This is very promising, as the signal in GW emission is proportional to the energy density in string loops. The asymptotic fraction of energy density in NS5-string loops shows an interesting dependence on the initial conditions of the field $\Phi_{0}$ and on the initial loop energy density fraction $\Omega^{(0)}_{\rm loop}$, as shown in Fig. 5.
@@
We then turned on a radiation background fluid in the system and repeated our analysis. Again, we performed a stability study of the fixed points and the results are shown in the map in Fig. 6. In a similar fashion to what was observed in the absence of radiation, the F-string attractor for $\lambda=3$ is the loop tracker Sánchez González et al. (2025). Nevertheless, depending on the initial fraction of energy density in radiation and in loops, the system spends quite a long time around the radiation tracker $\mathcal{S}$. This slows the field down and, in the absence of string loops, would prevent it from overshooting. However, if the loops are present, the system will eventually move towards $\mathcal{L}$, again leading to overshooting, unless the field reaches its minimum while the system is transitioning from $\mathcal{S}$ to $\mathcal{L}$. In this case, the loop energy density may still grow as shown in Fig. 8. Moreover, as argued above, as the field oscillates around the minimum, $\Omega_{\rm loops}/\Omega_{\Phi}$ stays constant. However, radiation redshifts faster than matter, and so the radiation density drops while $\Omega_{\Phi}$ and $\Omega_{\rm loop}$ grow. This leads to a growth in the energy density of cosmic string loops after $\Phi$ settles into its minimum, until it reaches its asymptotic value and stabilises there.
@@ Acknowledgments:
Appendix: Stability analysis
In this appendix we gather the analytic conditions for the linear stability of the various fixed points, summarized in Figs. 1 and 6, following the standard procedure outlined in e.g. Copeland et al. (1998).
.1 Loops without radiation
@@
  • $\mathcal{F}$: $0<\lambda<2$ and $\beta>(-2+\lambda^{2})/\lambda$
@@
.2 Loops with radiation
@@
  • $\mathcal{F}$: $0<\lambda<2$ and $\beta>\frac{-2+\lambda^{2}}{\lambda}$

### v2

@@ 4 Oscillating tension
$T_{\phi}\gg T_{R}\quad\Leftrightarrow\quad\epsilon m\ll 1\,$, (4.3)
one can approximate $a^{2}\dot{R}^{2}$ by its average $\langle a^{2}\dot{R}^{2}\rangle=1/2$, integrate (2.5) and reduce the dynamics to:
$\ddot{R}+\left(2H+\frac{\dot{\mu}}{2\mu}\right)\dot{R}\,+\frac{R}{\epsilon_{0}^{2}}\frac{\mu}{\mu_{0}}=0\,.$ (4.4)
@@
$\ddot{\sigma}+\left(\frac{\mu}{\mu_{0}\epsilon_{0}^{2}}+\frac{3}{16}\frac{\dot{\mu}^{2}}{\mu^{2}}-\frac{1}{4}\frac{\ddot{\mu}}{\mu}-\frac{\ddot{a}}{a}-\frac{1}{2}\frac{\dot{a}\dot{\mu}}{a\mu}\right)\sigma=0\,.$ (4.6)
Noting that when the modulus oscillates around its minimum, so does the string tension, (4.6) is a Hill equation. The oscillating tension may open up a rich dynamics for the circular strings, in particular the possibility of resonances.
As is often the analysis of Mathieu/Hill’s equations, it is convenient to rescale the time variable such that the period of the field oscillation is exactly $\pi$. We therefore define $\tau\equiv\frac{m}{2}t$ such that (4.6) becomes:
$\sigma^{\prime\prime}+\left(4\frac{\mu/\mu_{0}}{m^{2}\epsilon_{0}^{2}}+\frac{3}{16}\frac{\mu^{\prime 2}}{\mu^{2}}-\frac{1}{4}\frac{\mu^{\prime\prime}}{\mu}-\frac{a^{\prime\prime}}{a}-\frac{1}{2}\frac{a^{\prime}\mu^{\prime}}{a\mu}\right)\sigma=0\,.$ (4.7)
Equation (4.7) is an accurate description of the circular string dynamics, governed by (2.5) and (2.6), in the regime $m\epsilon_{0}\ll 1$. In this limit one may approximate (4.7) by keeping only the first term inside the brackets, allowing for an approximate WKB solution of the form:
$\sigma(\tau)=\sigma_{0}\left(\frac{\mu_{0}}{\mu(\tau)}\right)^{1/4}\cos\left[\frac{2}{m\epsilon_{0}}\int_{0}^{\tau}d\tilde{\tau}\left(\frac{\mu_{0}}{\mu(\tilde{\tau})}\right)^{1/2}+\delta\right],$ (4.8)
where $\delta$ is a phase fixed by the initial conditions. We see that the amplitude of the oscillations is modulated by the time variation of the tension. We stress that (4.8) holds both in flat space and in the presence of gravity. Note also that this result is independent of whether the variation of the string tension is monotonic (as in [20]) or oscillating, provided it varies sufficiently slowly, cf. (4.3). In the WKB regime the physical radius of the string scales as $R_{\rm max}\propto a^{-1}\mu^{-1/2}$, and so it grows if:
$\frac{d}{dt}(a\mu^{1/2})<0\Leftrightarrow\frac{\dot{\mu}}{\mu}<-2H\,$, (4.9)
that is, if the growth condition of [20] is respected.
4.1 Flat spacetime
In order to build some physical intuition for the behaviour of the strings with an oscillating tension, let us first analyse the flat space solutions of (2.5) and (2.6). In flat space $a={\rm const}$, $H=0$ and the scalar motion is harmonic as there is no friction:
$\phi=\phi_{0}\cos(mt)=\phi_{0}\cos(2\tau)\,.$ (4.10)
@@ 4.1 Flat spacetime
$\mu=\mu_{0}\ e^{-\sqrt{6}\beta\,\phi_{0}\cos(2\tau)}\,.$ (4.11)
From (4.5) and (4.8) one then sees that the radius of the string follows:
$R(\tau)\propto\mu(\tau)^{-1/2}\,$, (4.12)
@@
Having seen that in the WKB regime the system presents no instability, we adapt the standard methods of Floquet’s analysis (see e.g. [2]) to the current problem and look for instabilities when (4.3) is not satisfied. Defining the Floquet exponent $\mu_{F}\in\mathbb{R}$ via:
$R(\tau)=e^{\mu_{F}\,\tau}f(\tau)\,$, (4.13)
@@
Figure 13: No instability in flat space: growth exponent as a function of the timescale of the problem $\epsilon_{0}^{2}m^{2}$ (obtained from a numerical analysis of (2.5)-(2.6) with $\beta\phi_{0}=\sqrt{2/3}$.)
@@ 4.2 Curved spacetime
$H^{2}=\frac{1}{3M_{p}^{2}}\left(\frac{1}{2}\dot{\phi}^{2}+V(\phi)\right).$ (4.14)
As above, we are interested in a regime where the scalar is oscillating around the minimum of its potential. If the potential is well approximated by a quadratic term, the solutions to (4.2) and (4.14) are of the form:
@@
$\mu=\mu_{0}\ e^{-\sqrt{6}\beta\,\frac{\phi_{0}}{\tau}\cos(\tau)}\,.$ (4.16)
We see that in the late time limit, the variation of the tension becomes less significant. Taking into account this time dependence, one can rewrite the solution (4.8) in terms of the string radius as:
$R(\tau)\propto\frac{\mu(\tau)^{-1/2}}{a(\tau)}\,.$ (4.17)
@@ 5 Discussion and conclusions
The first part of this work has been devoted to studying whether the rolling volume mode actually settles in its minimum or overshoots it causing a decompactification limit. In fact, the typical LVS scalar potential for the canonically normalised volume mode $\Phi$ (3.2) features a minimum and a maximum. If $\Phi$ reaches the minimum with too much kinetic energy, the field will overshoot the maximum and run away to infinity. In the absence of any background fluid, this overshoot problem is present for generic initial conditions. The canonical way to overcome this problem is to introduce a fraction of radiation in the system. Depending on the specific embedding of the model, this solution may or may not be well justified. Therefore, following the work of [36], we studied how the inclusion of a string loop fluid in the dynamical system affects the overshooting problem.
@@
We first analysed the behaviour of the dynamical system including cosmic string loops in the absence of a radiation background. We performed a stability analysis of the fixed points, summarised in the stability map in Fig. 1. For an exponential potential with $\lambda=3$, as for the approximated LVS potential (3.19), we found that the final attractor of a system with only F-strings ($\beta=1/2$) is the loop tracker $\mathcal{L}$, confirming the result of [36]. For $\beta=1/2$, the kinetic energy of the field is too high to stop at the minimum, and the system overshoots. On the other hand, for effective strings coming from D3-branes wrapping an internal 2-cycle ($\beta=1/3$) or NS5-branes wrapping a 4-cycle ($\beta=1/6$), the final attractor of the system would be the mixed tracker $\mathcal{T}_{1}$. As the convergence towards this fixed point is slow, the system still spends a long time hovering around $\mathcal{L}$, even if this critical point is unstable. Hence, whether the field overshoots or converges to the minimum is still dependent on the fraction of kinetic energy at $\mathcal{L}$. For $\beta=1/3$, the situation is qualitatively similar to the F-string case, with a kinetic energy that is too high at the minimum to avoid overshooting. For $\beta=1/6$, the NS5-string loops are so effective at absorbing energy from the field that they can indeed prevent it from overshooting. In fact, for NS5-strings, at $\mathcal{L}$ the kinetic energy of the field constitutes only a small fraction of the total energy density, $\Omega_{\rm k}=\beta^{2}\simeq 0.03$, while the fraction of energy density in NS5-string loops gets over $97\%$, as shown in Fig. 5. Moreover, when the field starts oscillating around the minimum, the tension of the strings stops decreasing. Therefore, the loops start redshifting as matter, just like the field. This means that the ratio of their energy densities stays constant throughout the period of early matter domination, provided that the strings do not lose energy through GW or particle emission. The most interesting aspect is that this fraction is always rather high, ranging from $25\%$ to $60\%$. This is very promising, as the signal in GW emission is proportional to the energy density in string loops. The asymptotic fraction of energy density in NS5-string loops shows an interesting dependence on the initial conditions of the field $\Phi_{0}$ and on the initial loop energy density fraction $\Omega^{(0)}_{\rm loop}$, as shown in Fig. 6.
@@
We then turned on a radiation background fluid in the system and repeated our analysis. Again, we performed a stability study of the fixed points and the results are shown in the map in Fig. 7. In a similar fashion to what was observed in the absence of radiation, the F-string attractor for $\lambda=3$ is the loop tracker [36]. Nevertheless, depending on the initial fraction of energy density in radiation and in loops, the system spends quite a long time around the radiation tracker $\mathcal{S}$. This slows the field down and, in the absence of string loops, would prevent it from overshooting. However, if the loops are present, the system will eventually move towards $\mathcal{L}$, again leading to overshooting, unless the field reaches its minimum while the system is transitioning from $\mathcal{S}$ to $\mathcal{L}$. In this case, the loop energy density may still grow as shown in Fig. 9. Moreover, as argued above, as the field oscillates around the minimum, $\Omega_{\rm loops}/\Omega_{\Phi}$ stays constant. However, radiation redshifts faster than matter, and so the radiation density drops while $\Omega_{\Phi}$ and $\Omega_{\rm loop}$ grow. This leads to a growth in the energy density of cosmic string loops after $\Phi$ settles into its minimum, until it reaches its asymptotic value and stabilises there.
@@ Acknowledgments:
Appendix A Stability analysis
In this appendix we gather the analytic conditions for the linear stability of the various fixed points, summarized in Figs. 1 and 7, following the standard procedure outlined in e.g. [24].
A.1 Loops without radiation
@@
  • $\mathcal{M}$: $0<\lambda<2$ and $\beta>(-2+\lambda^{2})/\lambda$
@@
A.2 Loops with radiation
@@
  • $\mathcal{M}$: $0<\lambda<2$ and $\beta>\frac{-2+\lambda^{2}}{\lambda}$